# Practical Rateless Set Reconciliation

### v2

@@ 6. Implementation
Efficient incremental encoding. A key feature of Rateless IBLT is that it allows Alice to generate and send coded symbols one by one until Bob can decode. Suppose that Alice has generated coded symbols until index $i-1$, and now wishes to generate the $i$-th coded symbol. She needs to quickly find the source symbols that are mapped to it. A strawperson solution is to store alongside each source symbol the next index it is mapped to, and scan all the source symbols to find the ones mapped to $i$. However, this takes $O(|A|)$ time. In our implementation, we store pointers to source symbols in a heap. It implements a priority queue, where the priority is the index of the next coded symbol that a source symbol is mapped to. A smaller value indicates higher priority. This ensures that source symbols used for generating the next coded symbol are always at the head of the queue so that the encoder can access them efficiently without scanning all the source symbols.
@@ 7. Evaluation
Schemes compared. We compare with regular IBLT (Goodrich and Mitzenmacher, 2011; Eppstein et al., 2011), MET-IBLT (Lázaro and Matuz, 2023), PinSketch (Dodis et al., 2004), and Merkle tries (Yue et al., 2020). For Rateless IBLT, we use our implementation discussed in § 6. For regular IBLT and MET-IBLT, we implement each scheme in Python. We use the recommended parameters (Eppstein et al., 2011, § 6.1)(Lázaro and Matuz, 2023, §§ V-A, V-C), and allocate 8 bytes for the checksum and the count fields, respectively. For PinSketch, we use Minisketch (Naumenko et al., 2019, § 6), a state-of-the-art implementation (Wuille, [n. d.]) written in C++ and deployed in Bitcoin. For Merkle tries, we use the implementation in Geth (go-ethereum Authors, 2024a), the most popular client for Ethereum.
@@ 7.2. Computation Cost
The item size $\ell$ affects both encoding and decoding because it decides the time it takes to compute the XOR of two symbols, which dominates the computation cost in Rateless IBLT. Fig. 10 shows the relative slowdown when as $\ell$ grows from $8$ bytes to $32$ KB. Initially, the slowdown is sublinear (e.g., less than 4×4$\times$4 × when $\ell$ grows by 16×16$\times$16 × from $8$ to $128$ bytes) because the other costs that are independent of $\ell$ (e.g., generating the mappings) are better amortized. However, after $2$ KB, the slowdown becomes linear. This implies that the data rate at which the encoder can process source symbols, measured in bytes per second, stays constant. For example, when encoding for $d=1000$, the encoder can process source symbols at $124.8$ MB/s. The same analysis applies to decoding. In comparison, tæhe encoding complexity of PinSketch increases linearly with $\ell$, and the decoding complexity increases quadratically (Dodis et al., 2004; Wuille, [n. d.]).
@@ 7.3. Application
Ethereum (as well as most other blockchains) currently uses Merkle tries (§ 2) to synchronize ledger states between replicas. It applied a few optimizations: using a 16-ary trie instead of a binary one, and shortening sub-tries that have no branches. The protocol is called state heal and has been deployed in Geth (go-ethereum Authors, 2024a), the implementation which accounts for 84% of all Ethereum replicas (one-three-three seven, 2024). Variants of Geth also power other major blockchains, such as Binance Smart Chain and Optimism.
@@
Results. We first vary the state snapshot that Bob loads and measure the completion time and the communication cost for Bob to synchronize with Alice. We fix the bandwidth to 20 Mbps. Fig. 11 shows the results. As Bob’s state becomes more stale, more updates happen between his and the latest states, and the difference between the two grows linearly. As a result, the completion time and the communication cost of both schemes increase linearly. Meanwhile, Rateless IBLT consistently achieves 5.65.65.65.613.6×13.6$\times$13.6 × lower completion time, and 4.44.44.44.46.9×6.9$\times$6.9 × lower communication cost compared to state heal. As discussed previously, state heal has a much higher communication cost because it requires transmitting the differing internal nodes of the Merkle trie in addition to the leaves. For example, this amplifies the number of trie nodes transmitted by 3.6×3.6$\times$3.6 × when Bob’s state is 30 hours stale. The higher communication cost leads to proportionately longer completion time as the system is throughput-bound. We observe similar gains when Bob loads fresher states, e.g., ones less than 20 minutes stale, as shown in Fig. 15 in Appendix E.
In our experiments, state heal requires at least 11 rounds of interactivity, as Alice and Bob descend from the roots of their tries to the differing leaves in lock steps. Rateless IBLT, in comparison, only requires half of a round because Alice streams coded symbols without waiting for any feedback. This advantage is the most obvious when reconciling a small difference, where the system would be latency-bound. For example, Rateless IBLT is 8.2×8.2$\times$8.2 × faster than state heal when Bob’s ledger state is only 1 block (12 seconds) stale. See Fig. 16 in Appendix E for traces of this experiment that highlight the impact of interactivity.
@@ 8. Conclusion
Acknowledgements
We thank Francisco Lázaro for fruitful discussions. Lei Yang is supported by a gift from the Ethereum Foundation.
Appendix A Analysis on regular IBLTs
@@
Appendix B Derivation of $P_{g}$ and $C(x)$
Here we calculate $P_{g}$ and $C(x)$ as defined in § 4.2 when the parameter $\alpha$ in $\rho(i)$ is set to $0.5$.
@@ Appendix B Derivation of $P_{g}$ and $C(x)$
$=\frac{2}{i+g+2}\prod_{n=1}^{g-1}\frac{i+n}{i+n+2}$
$=\frac{2}{i+g+2}\left(\prod_{n=1}^{g-1}i+n\right)\left(\prod_{n=3}^{g+1}\frac{1}{i+n}\right)$
$=\frac{2(i+1)(i+2)}{(i+g)(i+g+1)(i+g+2)}.$
Before calculating $C(x)$, notice that
$P_{g}=\frac{(i+1)(i+2)}{(i+g)(i+g+1)}-\frac{(i+1)(i+2)}{(i+g+1)(i+g+2)}.$
@@
$=\sum_{g=1}^{x}\left(\frac{(i+1)(i+2)}{(i+g)(i+g+1)}-\frac{(i+1)(i+2)}{(i+g+1)(i+g+2)}\right)$
$=\sum_{g=1}^{x}\frac{(i+1)(i+2)}{(i+g)(i+g+1)}-\sum_{g=2}^{x+1}\frac{(i+1)(i+2)}{(i+g)(i+g+1)}$
$=1-\frac{(i+1)(i+2)}{(i+x+1)(i+x+2)}$
$=\frac{x(2i+x+3)}{(i+x+1)(i+x+2)}.$
@@ Appendix C Deferred Proofs
[…]
@@
Next, let us consider $\Lambda(x)$. For a random source symbol, it is mapped to the $i$-th coded symbol independently with probability $\rho(i)$. Its degree $\Lambda$ is thus the sum of independent Bernoulli random variables with success probabilities $\rho(0),\rho(1),\dots,\rho(m-1)$, which follows Poisson binomial distribution. By an extension (Steele, 1994, § 5) to Le Cam’s theorem (Le Cam, 1960), we can approximate this distribution with a Poisson distribution of rate $\sum_{i=0}^{m-1}\rho(i)$, i.e., $\mathbb{E}(\Lambda)$, with the total variation distance between the two distributions tending to zero as $m$ goes to infinity. That is,
@@
Appendix D Additional Analytical Results
Theorem 5.1 shows the relationship between $\alpha$, the parameter in the mapping probability function $\rho(i)=\frac{1}{1+\alpha i}$, and the communication overhead. For any $\alpha$, the overhead $\eta^{*}$ is the smallest $\eta$ that satisfies Eq. 2. Fig. 14 shows the relationship between $\alpha$ and $\eta^{*}$.
Figure 14. Relationship between the communication overhead and the parameter $\alpha$.
Appendix E Additional Evaluation Results
Figure 15. Completion time and communication cost when synchronizing Ethereum ledger states at different staleness over a network link with 50 ms of propagation delay and 20 Mbps of bandwidth.
Fig. 15 complements Fig. 11 (§ 7.3). It shows the completion time and communication cost when synchronizing a smaller ledger state difference. Rateless IBLT achieves 4.84.84.84.810.9×10.9$\times$10.9 × lower completion time than state heal, and 6.96.96.96.98.6×8.6$\times$8.6 × lower communication cost.
Fig. 16 shows traces of bandwidth usage when synchronizing one block worth of state difference (see § 7.3 for the setup). For Rateless IBLT, the first coded symbol arrives at Bob in 1 round-trip time (RTT) after his TCP socket opens (0.5 RTT for TCP ACK to reach Alice, and another 0.5 RTT for the first symbol to arrive). Subsequent symbols arrive at line rate, as the peak at 1 RTT indicates. In comparison, for state heal, Alice and Bob reach the bottom of their tries after 11 RTTs; before that, they do not know the actual key-value pairs that differ, and the network link stays almost idle.

### v3

@@ 6. Implementation
Efficient incremental encoding. A key feature of Rateless IBLT is that it allows Alice to generate and send coded symbols one by one until Bob can decode. Suppose that Alice has generated coded symbols until index $i-1$, and now wishes to generate the $i$-th coded symbol. She needs to quickly find the source symbols that are mapped to it. A strawman solution is to store alongside each source symbol the next index it is mapped to, and scan all the source symbols to find the ones mapped to $i$. However, this takes $O(|A|)$ time. In our implementation, we store pointers to source symbols in a heap. It implements a priority queue, where the priority is the index of the next coded symbol that a source symbol is mapped to. A smaller value indicates higher priority. This ensures that source symbols used for generating the next coded symbol are always at the head of the queue so that the encoder can access them efficiently without scanning all the source symbols.
@@ 7. Evaluation
Schemes compared. We compare with regular IBLT (Goodrich and Mitzenmacher, 2011; Eppstein et al., 2011), MET-IBLT (Lázaro and Matuz, 2023), PinSketch (Dodis et al., 2008), and Merkle tries (Yue et al., 2020). For Rateless IBLT, we use our implementation discussed in § 6. For regular IBLT and MET-IBLT, we implement each scheme in Python. We use the recommended parameters (Eppstein et al., 2011, § 6.1)(Lázaro and Matuz, 2023, §§ V-A, V-C), and allocate 8 bytes for the checksum and the count fields, respectively. For PinSketch, we use Minisketch (Naumenko et al., 2019, § 6), a state-of-the-art implementation (Wuille, 2018) written in C++ and deployed in Bitcoin. For Merkle tries, we use the implementation in Geth (go-ethereum Authors, 2024a), the most popular client for Ethereum.
@@ 7.2. Computation Cost
The item size $\ell$ affects both encoding and decoding because it decides the time it takes to compute the XOR of two symbols, which dominates the computation cost in Rateless IBLT. Fig. 11 shows the relative slowdown when as $\ell$ grows from $8$ bytes to $32$ KB. Initially, the slowdown is sublinear (e.g., less than 4×4$\times$4 × when $\ell$ grows by 16×16$\times$16 × from $8$ to $128$ bytes) because the other costs that are independent of $\ell$ (e.g., generating the mappings) are better amortized. However, after $2$ KB, the slowdown becomes linear. This implies that the data rate at which the encoder can process source symbols, measured in bytes per second, stays constant. For example, when encoding for $d=1000$, the encoder can process source symbols at $124.8$ MB/s. The same analysis applies to decoding. In comparison, the encoding complexity of PinSketch increases linearly with $\ell$, and the decoding complexity increases quadratically (Dodis et al., 2008; Wuille, 2018).
@@ 7.3. Application
Ethereum (as well as most other blockchains) currently uses Merkle tries (§ 2) to synchronize ledger states between replicas. It applied a few optimizations: using a 16-ary trie instead of a binary one, and shortening sub-tries that have no branches. The protocol is called state heal and has been deployed in Geth (go-ethereum Authors, 2024a), the implementation which accounts for 84% of all Ethereum replicas (Sonic, 2024). Variants of Geth also power other major blockchains, such as Binance Smart Chain and Optimism.
@@
Results. We first vary the state snapshot that Bob loads and measure the completion time and the communication cost for Bob to synchronize with Alice. We fix the bandwidth to 20 Mbps. Fig. 12 shows the results. As Bob’s state becomes more stale, more updates happen between his and the latest states, and the difference between the two grows linearly. As a result, the completion time and the communication cost of both schemes increase linearly. Meanwhile, Rateless IBLT consistently achieves 4.84.84.84.813.6×13.6$\times$13.6 × lower completion time, and 4.44.44.44.48.6×8.6$\times$8.6 × lower communication cost compared to state heal. As discussed previously, state heal has a much higher communication cost because it requires transmitting the differing internal nodes of the Merkle trie in addition to the leaves. For example, this amplifies the number of trie nodes transmitted by 3.6×3.6$\times$3.6 × when Bob’s state is 30 hours stale. The higher communication cost leads to proportionately longer completion time as the system is throughput-bound.
(a) Staleness between 20 minutes and 100 hours.
(b) Staleness between 1 minute and 20 minutes.
@@
In our experiments, state heal requires at least 11 rounds of interactivity, as Alice and Bob descend from the roots of their tries to the differing leaves in lock steps. Rateless IBLT, in comparison, only requires half of a round because Alice streams coded symbols without waiting for any feedback. This advantage is the most obvious when reconciling a small difference, where the system would be latency-bound. For example, Rateless IBLT is 8.2×8.2$\times$8.2 × faster than state heal when Bob’s ledger state is only 1 block (12 seconds) stale.
We quickly highlight the impact of interactivity. Fig. 13 shows traces of bandwidth usage when synchronizing one block worth of state difference. For Rateless IBLT, the first coded symbol arrives at Bob in 1 round-trip time (RTT) after his TCP socket opens (0.5 RTT for TCP ACK to reach Alice, and another 0.5 RTT for the first symbol to arrive). Subsequent symbols arrive at line rate, as the peak at 1 RTT indicates. In comparison, for state heal, Alice and Bob reach the bottom of their tries after 11 RTTs; before that, they do not know the actual key-value pairs that differ, and the network link stays almost idle.
@@
8. Irregular Rateless IBLTs
When designing Rateless IBLTs (§ 4.1), the key task was to define the mapping rule that decides whether a source symbol $x$ should be mapped to the $i$-th coded symbol. Lemmas 4.1 and 4.2 showed that the mapping rule cannot be uniform over coded symbols: the probability that a random source symbol is mapped to the $i$-th coded symbol must decrease with $i$. In other words, different coded symbols are statistically inequivalent. On average, a coded symbol with a smaller index sees more source symbols mapped to it than one with a larger index does.
However, the same is not true for source symbols in our design: every subset of source symbols uses the same mapping probability $\rho(i)=\frac{1}{1+\alpha i}$ with the same parameter $\alpha$. This leaves a degree of freedom which we did not explore. We may divide source symbols into multiple subsets and use different $\rho(i)$ (in particular, different $\alpha$) for each subset. Similar techniques have successfully improved the communication costs of regular IBLTs (Lázaro and Matuz, 2021). In this section, we apply this technique on Rateless IBLTs and discuss the implications.
Concretely, we partition source symbols into $c$ mutually exclusive subsets, where $c$ is a parameter. A random source symbol belongs to subset $j$ ($0\leq j<c$) with probability $w_{j}$, which is another set of parameters. For a source symbol $x$, we choose the subset it belongs to based on its hash. For example, $x$ belongs to subset $j$ if $\sum_{b=0}^{j-1}w_{b}\leq\mathtt{Hash}(x)<\sum_{b=0}^{j}w_{b}$ assuming the hash is uniformly distributed in $[0,1)$. For each subset $j$, we define a parameter $\alpha_{j}$ and use mapping probability $\rho_{j}(i)=\frac{1}{1+\alpha_{j}i}$ when mapping source symbols in this subset. In other words, we replace $\alpha$ with a subset-specific $\alpha_{j}$ in the algorithm described in § 4.2. By convention, we call this generalized design Irregular Rateless IBLTs. Rateless IBLTs as discussed prior to this section is a special case where $c=1$, $w_{0}=1$, and $\alpha_{0}=0.5$.
As mentioned, the main benefit of Irregular Rateless IBLTs over Rateless IBLTs is a lower communication cost. Unfortunately, the density evolution analysis does not produce a closed-form result like the one in Theorem 5.1. To find a good configuration of $c$, $w_{j}$, and $\alpha_{j}$ that minimizes the overhead, we use brute force and try different values in simulations. To limit the complexity of the search, we set the number of subsets $c$ to $3$ and found the following optimal configuration
$c=3,$
$w_{0},w_{1},w_{2}=0.18,0.56,0.26,$
$\alpha_{0},\alpha_{1},\alpha_{2}=0.11,0.68,0.82.$
As shown in Fig. 15, the resulting communication overhead converges to $1.10$, which is $19\%$ lower than Rateless IBLTs (§ 4) and only $10\%$ above the information-theoretic lower bound. Meanwhile, encoding and decoding are $1.88$ times slower than Rateless IBLTs. As mentioned in § 4.2, the main reason is that computing mappings when $\alpha\neq 0.5$ requires raising numbers to arbitrary non-integer powers, while the case of $\alpha=0.5$ only requires computing square roots, which is faster on modern hardware. We leave further optimizations of the parameters and the implementation to future works.
Figure 15. Communication overhead of Rateless IBLTs and Irregular Rateless IBLTs as the difference size changes. We run 100 simulations for each data point and report the average.
@@ 9. Conclusion
We point out a few interesting future directions: optimizing the parameters and the implementation of Irregular Rateless IBLTs; considering scenarios where Alice and Bob’s sets change in the middle of reconciliation; and designing efficient solutions for reconciliation across more than two parties.
Acknowledgments
We thank Francisco Lázaro for fruitful discussions. Lei Yang was supported by a gift from the Ethereum Foundation. Yossi Gilad was partially supported by the Alon Fellowship.
This work does not raise any ethical issues. Appendices are supporting material that has not been peer-reviewed.
Appendix A Inflexibility of Regular IBLTs
@@
Appendix B Calculation of $P_{g}$ and $C(x)$
In this section, we calculate $P_{g}$ and $C(x)$ as defined in § 4.2.
@@ Appendix B Calculation of $P_{g}$ and $C(x)$
$=\frac{1/\alpha}{i+g+1/\alpha}\prod_{n=1}^{g-1}\frac{i+n}{i+n+1/\alpha}$
$=\frac{(i+1)_{g-1}}{\alpha(i+1+1/\alpha)_{g}}.$
Here, $(x)_{n}$ is the Pochhammer symbol.888$(x)_{n}=x(x+1)\dots(x+n-1)$.
Before proceeding to calculate $C(x)$, we first prove a useful identity about quotients of Gamma functions,
$\frac{\Gamma(x)}{\Gamma(x+y)}\equiv\frac{1}{y-1}\left(\frac{\Gamma(x)}{\Gamma(x+y-1)}-\frac{\Gamma(x+1)}{\Gamma(x+y)}\right).$
We start from the right-hand side,
$\frac{1}{y-1}\left(\frac{\Gamma(x)}{\Gamma(x+y-1)}-\frac{\Gamma(x+1)}{\Gamma(x+y)}\right)$
$=$ $\frac{1}{y-1}\left(\frac{(x+y-1)\Gamma(x)}{(x+y-1)\Gamma(x+y-1)}-\frac{\Gamma(x+1)}{\Gamma(x+y)}\right)$
$=$ $\frac{1}{y-1}\left(\frac{(y-1)\Gamma(x)+\Gamma(x+1)}{\Gamma(x+y)}-\frac{\Gamma(x+1)}{\Gamma(x+y)}\right)$
$=$ $\frac{\Gamma(x)}{\Gamma(x+y)}.$
The identity immediately implies the following
(4) $\sum_{x=a}^{b}\frac{\Gamma(x)}{\Gamma(x+y)}\equiv\frac{1}{y-1}\left(\frac{\Gamma(a)}{\Gamma(a+y-1)}-\frac{\Gamma(b+1)}{\Gamma(b+y)}\right).$
@@
$=\sum_{g=1}^{x}\frac{(i+1)_{g-1}}{\alpha(i+1+1/\alpha)_{g}}$
$=\sum_{g=1}^{x}\frac{\Gamma(i+g)\Gamma(i+1+1/\alpha)}{\alpha\Gamma(i+1)\Gamma(i+1+g+1/\alpha)}$
$=\frac{\Gamma(i+1+1/\alpha)}{\alpha\Gamma(i+1)}\sum_{g=1}^{x}\frac{\Gamma(i+g)}{\Gamma(i+g+1+1/\alpha)}$
$=\frac{\Gamma(i+1+\frac{1}{\alpha})}{\Gamma(i+1)}\left(\frac{\Gamma(i+1)}{\Gamma(i+1+\frac{1}{\alpha})}-\frac{\Gamma(i+x+1)}{\Gamma(i+x+1+\frac{1}{\alpha})}\right)$
$=1-\frac{\Gamma(i+1+\frac{1}{\alpha})\Gamma(x+i+1)}{\Gamma(i+1)\Gamma(x+i+1+\frac{1}{\alpha})}.$
The second last equality results from applying Eq. 4.
@@ Appendix C Deferred Proofs
$\forall q\in(0,1]:e^{\frac{1}{\alpha}\text{Ei}\left(-\frac{q}{\alpha\eta}\right)}<q.$
@@
Next, let us consider $\Lambda(x)$. For a random source symbol, it is mapped to the $i$-th coded symbol independently with probability $\rho(i)$. Its degree $\Lambda$ is thus the sum of independent Bernoulli random variables with success probabilities $\rho(0),\rho(1),\dots,\rho(m-1)$, which follows Poisson binomial distribution. By an extension (Steele, 1994, § 5) to Le Cam’s theorem (Cam, 1960), we can approximate this distribution with a Poisson distribution of rate $\sum_{i=0}^{m-1}\rho(i)$, i.e., $\mathbb{E}(\Lambda)$, with the total variation distance between the two distributions tending to zero as $m$ goes to infinity. That is,